### v1

# A Holistic Optimization Framework for Energy Efficient UAV-assisted Fog Computing: Attitude Control, Trajectory Planning and Task Assignment

Abstract
Unmanned Aerial Vehicles (UAVs) have significantly enhanced fog computing by acting as both flexible computation platforms and communication mobile relays. In this paper, we propose a holistic framework that jointly optimizes the total latency and energy consumption for UAV-assisted fog computing in a three-dimensional spatial domain with varying terrain elevations and dynamic task generations. Our proposed framework considers three important and interdependent modules: attitude control, trajectory planning, and task assignment. We first establish a fuzzy proportional-integral-derivative control model to determine the UAV’s attitude. Then, we propose an enhanced Ant Colony System (ACS) based algorithm, that includes a safety value and a decoupling mechanism to overcome the convergence issue in classical ACS, to compute the optimal UAV trajectory. Finally, we design an algorithm based on the Particle Swarm Optimization technique, to determine where each offloaded task should be executed. Under our proposed framework, the outcome of one module would affect the decision-making in one other, providing a holistic perspective of the system and thus leading to improved solutions. We demonstrate by extensive simulation results that our proposed framework can significantly improve the overall performance, measured by latency and energy consumption, compared to existing baseline approaches.
Index Terms:
Fog Computing, Unmanned Aerial Vehicles (UAV), Attitude Control, Trajectory Planning, Ant Colony Algorithm
I Introduction
Driven by the development of the Internet of Things (IoT) [1], mobile terminal devices such as smartphones and tablets are now capable of generating and collecting massive amounts of data. However, the ability of processing these data, such as performing computational tasks, in the IoT devices are still limited [2, 3]. On the other hand, Unmanned Aerial Vehicles (UAVs) have been recently identified as a versatile platform that connects IoT devices and servers or data centers via the network edge [4]. In addition, some UAVs are equipped with computational capabilities and thus can be regarded as “moving fog nodes” for offloading certain computational tasks. To fully utilize the versatility and flexibility of UAVs in fog computing, key considerations include the management of each UAV’s attitude, the strategic planning of their respective trajectories, and the efficient task assignment policy to determine the appropriate computing device for each task.
This paper considers a fog computing environment with a single UAVs deployed at the network edge, aiming at maximizing the energy efficiency by collaboratively controlling the attitudes and planning the trajectories, as well as determining the location for each computing task to be executed. Without loss of generality, we focus on quadrotor UAVs, which are considered ideal for computation offloading in IoT networks due to their enhanced flexibility [5, 6, 7, 8].
Attitude control, trajectory planning, and task assignment policy in UAV-assisted fog computing are very important processes that significantly impact the overall efficiency of the system [9, 10]. Specifically, attitude control (by adjusting pitch, roll, and yaw) ensures a stable and precise orientation of the UAV during operation, which is a necessary condition to maintain a high quality of communication with IoT devices and other fog nodes [11]. In addition, a stable attitude facilitates the UAV to effectively perform computation and storage tasks as a fog node. On the other hand, trajectory planning of the UAV can reduce its power consumption by identifying the most efficient path to collect data and tasks based on the locations of IoT devices [12]. Finally, task assignment refers to the process of deciding whether a specific task should be handled locally by the IoT device, processed in the fog layer (including the UAV), or offloaded to the central cloud. Specific objectives, including minimizing latency, maximizing throughput, or optimizing energy efficiency, can be achieved by smartly assigning tasks to appropriate devices. Task assignment is often jointly optimized with resource allocation, where transmission resources such as power and bandwidth are distributed among different transmission pairs in the network, to facilitate the transmission process and improve the overall efficiency. The four processes are inherently linked in UAV-assisted fog computing. For example, a UAV’s attitude control would ensure that it maintains optimal orientations while following a planned trajectory or processing a task. Also, when deciding whether to offload a certain task to the fog or the cloud and how to allocate relevant transmission resources, the energy consumption and latency for a certain UAV to reach the proximity of the IoT initiating the task along a planned trajectory should also be taken into account.
Existing studies have considered two or three processes for joint optimization. For example, Wei et al. [13] proposed three decision-making algorithms to solve the joint optimization problem involving energy consumption and mean delay. Our previous published conference paper [14] considered the joint optimization of trajectory planning and task assignment, in a fog computing system assisted by one single UAV. We proposed an optimization framework that integrated the classical Ant Colony System (ACS, also referred as Ant Colony Optimization/ACO in some existing studies) and Particle Swarm Optimization (PSO) algorithms. The framework would generate the optimal trajectory in a two-dimensional plane with randomly generated computing tasks and obstacles, and decide the optimal location (the IoT device, the UAV, or the central cloud) for each task to be executed.
In this work, we revise and extend the model in [14] to better align with the practical scenarios of UAV-assisted fog computing. Particularly, we consider a three-dimensional network space (compared to a two-dimensional plane as in [14]) with varying terrain elevations. We also add the attitude control and optimal structure design modules for a quadrotor UAV, to ensure the stability of flights, prevent potential structural damage, and reduce the latency when the UAV ascends or descends by two newly proposed Active Disturbance Rejection Control (ADRC) strategies [15]. Compared to most existing similar studies, (e.g.,[16, 6, 7, 8]), which only considered the flight attitude at independent static points, we jointly consider planning the optimal trajectory and determining the attitudes, taking into account the extra consumption required for changing attitude along the trajectory. Our research is expected to provide new useful insights in UAV-assisted fog computing applications for improving the overall performance and efficiency.
The contributions of this paper are summarized as follows.
  • We propose a holistic framework with multiple modules including attitude control, trajectory planning, resource allocation, and task assignment, to optimize the overall energy efficiency for UAV-assisted fog computing in a three-dimensional spatial domain with varying terrain elevations. To the best of our knowledge, this is the first work that jointly considers these modules together and takes advantage of their intricate interconnections. By integrating and coordinating the modules, our proposed framework is capable of achieving optimal performance and energy efficiency in the dynamic and complex environment of UAV-assisted fog computing.
  • We establish a fuzzy proportional-integral-derivative control (PID) model for quadrator UAV attitude control. Compared to the classical PID control commonly used in existing studies, the fuzzy PID control can capture the dependencies among parameters and adapt to the dynamic environment. We demonstrate that the fuzzy PID control can effectively improve the stability in takeoff, cruise, and landing phases, and thus reduce both latency and energy consumption in UAV-assisted fog computing.
  • We propose a computationally efficient algorithm called ACS-DS (ACS with Decoupling and Safety values), to incorporate two mechanisms for decoupling and safety values in the ACS algorithm, to overcome the common issues of slow convergence and trapped in local optimum in classical ACS in UAV trajectory planning problems. We prove the convergence of the proposed algorithm and show that it has a polynomial time complexity. We also verify its superiority in terms of the convergence performance over classical ACS by numerical experiments.
  • We propose a heuristic algorithm based on the idea of PSO to solve the joint optimization problem involving resource allocation and task assignment in UAV-assisted fog computing, given that an initial trajectory of the UAV has been determined. The heuristic algorithm is shown to be capable of overcoming the complexities in the non-convex optimization problem and obtaining the optimal task assignment and resource allocation decisions efficiently.
  • We demonstrate, through extensive numerical experiments, that our proposed holistic framework can reduce overall operational efficiency cost by more than $58\%$ compared to baseline implementations.
The rest of this paper is organized as follows. Section II reviews recent advancements on attitude control, trajectory planning, and task assignment in UAV-assisted fog computing. Section III provides descriptions on the UAV architecture as well as key metrics at the system level. Section IV explains the formulation of the joint optimization problem. Section V describes the proposed computationally efficient algorithms to solve the problem in detail. Section VI demonstrates the improvements of the proposed algorithm by extensive numerical results. Section VII concludes the paper.
II Related Work
The fuzzy proportional-integral-derivative (PID) control system was adopted in attitude control and trajectory planning processes [17, 18]. It has been demonstrated that a proper combination of proportional-integral (PI) and proportional-derivative (PD) values and an optimally tuned fuzzy gain can better stabilize and control the attitude of quadrotor UAVs, as well as track their trajectories with smaller errors.
For trajectory planning of UAVs, heuristic algorithms such as PSO [19], ACS [20], and genetic algorithm (GA) [21], have been adopted to overcome the space and computation complexities in such problems. Compared with another branch of approaches that use neural networks and deep learning as the key techniques (e.g., [22, 23, 24]), heuristic algorithms are more interpretable and less data dependent, and thus more appropriate for UAV-assisted fog computing scenarios where the operational environments are usually highly diverse and dynamic, and transparent heuristics are preferred for regulators to validate and verify the operation. Among the heuristic algorithms, the PSO [25] is one of the most commonly used techniques in such problems. However, one major concern of applying PSO in complex systems is that PSO may be converged prematurely to local optima [26].
ACS, based on the study of ants searching for food, is another commonly adopted technique in discrete and continuous optimization problems [27, 28]. While ACS-based approaches are well-known for their robustness, they also suffer the disadvantage of being prone to local optima due to premature convergence like PSO. To overcome this issue, Ali et al. proposed the maximum-minimum ACS that improved the performance at the expense of slower convergence speed [29]. In addition, Wang et al. [30] presented an adaptive double-layer ant colony optimization algorithm (DL-ACS) based on an elitist strategy (ADAS) and an improved moving average algorithm (IMA) to solve a three-dimensional UAV trajectory planning problem. Liu et al. [31] also proposed an efficient online path-planning algorithm based on DL-ACS, that was shown to improve the optimal solution and convergence speed concurrently, and to enhance the stability of the algorithm.
Meanwhile, resource allocation or task assignment problems have been extensively studied in existing studies. For example, Cui et al. [32] proposed a multi-agent Q-learning-based reinforcement learning (MARL) framework, where each agent independently executes the allocation algorithm to optimize the overall energy efficiency. Li et al. [33] considered a joint optimization problem involving trajectory optimization and task allocation, with the goal of minimizing UAV energy consumption and optimizing computation offloading and using successive convex approximation (SCA) technique to solve it. Wu et al. [34] proposed a cooperative multi-agent deep reinforcement learning framework, which combines task assignment and allocation of limited communication resources to minimize the overall energy consumption and delay.
In trajectory planning and/or task assignment problems, one notable issue of applying ACS or PSO is that the algorithm may enter the deadlock state where one or more tasks wait endlessly for resources [35]. Zhou et al. [36] and Hou et al. [37] both proposed ACS-based algorithms with enhanced communication mechanisms to avoid deadlocks. However, while the proposed methods managed to reduce the likelihood of deadlocks, they did not eliminate the possibilities of such undesirable events. Another effort to overcome the deadlock problem is to establish a dead-corner table (or referred as a dead angle table[38, 39], which updates the path strength using a penalty function to avoid the deadlock. While this approach effectively overcomes the deadlock problem, it introduces additional overheads in maintaining the table and thus is less efficient.
TABLE I: A summary and comparison of related studies
Research Process(es) focused Method(s) Objective(s)
[32] Trajectory Planning & Task Assignment MARL Max. long-term rewards
[33] Trajectory Planning & Task Assignment SCA Max. UAV energy efficiency
[19] Task Assignment PSO Min. the overall workflow execution cost
[21] Task Assignment GA Min. communication costs
[13] Task Assignment Reinforcement learning Max. the platform’s utility
[29] Trajectory Planning MMACS Min. route length while avoiding collision
[26] Trajectory Planning Multi-Swarm PSO Max. the survivability of UAV
[31] Trajectory Planning Double-layer ACS Max. the distribution efficiency
[30] 3D-Trajectory Planning Double-layer ACS Max. Allocation Efficiency
[31] Trajectory Planning Double-layer ACS Max. Allocation Efficiency
[36],[37] Trajectory Planning Enhanced ACS Avoid path deadlock
[38],[39] Trajectory Planning Enhanced ACS & Dead-corner table Avoid path deadlock
[17] Attitude Control Traditional PID controller Control attitude
[18] Trajectory Planning & Attitude Control Fuzzy PID controller Control attitude and track trajectory
[34] Task Assignment Reinforcement learning Min. energy consumption and latency
Our Work 3D Trajectory Planning & Task Assignment & Attitude Control PSO & Fuzzy control & ACS-DS Min. energy consumption and latency
Our work in this paper combines the techniques mentioned above that have been demonstrated to be effective in respective processes, including fuzzy PID attitude control, ACS-based trajectory planning, and PSO-based task allocation and resource allocation. Following this foundation, we further improve several aspects of the entire framework, such as 1) the hardware of the quadrotor is designed to match the efficient performance of the fuzzy PID control system, 2) two enhanced anti-lockout mechanisms (decoupling and safety values) in the classical ACS for trajectory planning, and 3) a modified PSO approach with improved efficiency that is more appropriate for task assignment and resource allocation in large-scale UAV-assisted fog computing systems.
III System Model
III-A Network Structure and Components
We consider a UAV-assisted fog computing system in a three-dimensional Euclidean space. The system consists of a quadrotor UAV, a remote data center (DC) in the cloud, and $K$ mobile IoT devices (MDs). A demonstration of the key structures in the system is shown in Fig. 1.
Figure 1: The structure of UAV-assisted fog computing network.
The positions of $K$ MDs are randomly generated according to a Poisson Point Process (PPP), with the coordinate of the $j$th MD denoted by $\boldsymbol{M}_{j}=\left[x_{j},y_{j},z_{j}\right]$. For convenience, we discretize a continuous time horizon with length $T$ uniformly into $N$ timeslots, and thus each timeslot has a length of $L=\frac{T}{N}$. We assume that $N$ is sufficiently large, or equivalently, the timeslots are sufficiently short, such that the position of the UAV can be considered to be fixed within each timeslot. We use $[x(t),y(t),z(t)]$, where $t\in\{1,2,\cdots,N\}$, to denote the position of the UAV at the $t$th timeslot. The UAV is powered by a battery with a maximum capacity of $B_{c}$. The height and speed are restricted to not exceed $z^{\max}$ and $v^{\max}$, respectively, at all times.
We consider that task arrivals from each MD conform to a Poisson process, with an arrival rate $\lambda_{j}$ from the $j$th MD. We use $s_{ij}(t)$, which is assumed to be exponentially distributed, to denote the data size of the $i$th task from the $j$th MD to be processed at the $j$th timeslot, and $c_{ij}$ to denote the number of CPU cycles required to process the task.
As we introduced in [14], the proportion of channels allocated to the $j$th MD would be based on the Gamma distribution,
$P_{j}(t)=\frac{\beta^{\alpha}}{(\alpha-1)!}s_{ij}(t)^{\alpha-1}e^{-\beta s_{ij}(t)},$ (1)
where $\alpha$ and $\beta$ are the shape and rate parameters in the Gamma distribution. For demonstration purposes, we set $\alpha=\beta=2$ throughout the paper. Given the total number of available channels $N_{c}$, the number of channels allocated to the $j$th MD at the $t$th timeslot is $C_{j}(t)=\left[N_{c}\cdot P_{j}(t)\right]$ where $\left[x\right]$ rounds $x$ to the nearest integer. The rationale of this arrangement is to prevent the channel from being monopolized by extremely large tasks, and thus to ensure a certain level of transmission efficiency for all tasks.
For the $i$th task from the $j$th MD, we denote the energy consumption and delay as $E_{ij}$ and $D_{ij}$, respectively. For each task, the quadrotor UAV needs to move sufficiently close to the MD that initiates the task through a planned trajectory, in order to receive and further process the task. As mentioned earlier, a task may be executed locally, in the fog layer by the UAV, or further offloaded to the data centers in the central cloud. The computing results of any offloaded task need to be transmitted back to the initiating MD. In this context, we define $x_{ij}(t)=(x_{ij}^{\text{MD}}(t),x_{ij}^{\text{UAV}}(t),x_{ij}^{\text{DC}}(t))$ as an array consisting of binary variables that indicate the specific execution location of the $i$th task from the $j$th MD during the $t$th timeslot. For ease of reference, we list the key notations related to the system model in Table II.
TABLE II: Key notations in the system model
Notation Definition
$K$ The total number of MDs
$L$ The length of every timeslot
$T$ The total number of timeslots required to complete all tasks
$(x_{j},y_{j},z_{j})$ The location of the $j$th MD
$(x(t),y(t),z(t))$ The location of the UAV at the $t$th timeslot
$v(t)$ The UAV’s velocity vector at the $t$th timeslot
$d_{j}(t)$ The distance between the UAV and $j$th MD at the $t$th timeslot
$B_{c}$ The battery capacity of the UAV
$N_{j}$ The total number of tasks from the $j$th MD
$N_{c}$ The total number of wireless channels
$C_{j}(t)$ The number of channels assigned to the $j$th MD at the $t$th timeslot
$N_{b}$ The number of UAV propeller blades
$R$ The radius of UAV propeller blades
$\gamma$ The mounting angle of UAV propeller blades
$P_{w},P_{t}$ Width and thickness of UAV propeller blades
$p_{u}(t)$, $p_{D}(t)$, $p_{j}(t)$ The transmission power of the UAV, the DC, or the $j$th MD at the $t$th timeslot
$f(t)$, $f_{j}(t)$ The processing frequency of equipment for UAV, or the $j$th MD at the $t$th timeslot
$\mathbb{S}(t)$ The weighted sum of network energy consumption and delay at the $t$th timeslot
III-B The Controllable Structure of Quadrotor UAV
Quadrotor UAVs are capable of achieving six degrees of freedom in pitch, yaw, roll, vertical, fore and aft, and lateral motion by translating and rotating along the $x$, $y$ and $z$ axes. The process is accomplished by controlling four kinetic input quantities of the motor, denoted by $\mathbf{U}=\left(U_{1},U_{2},U_{3},U_{4}\right)^{T}$, and six output control variables that include the changes of movements $\mathbf{V}=(\Delta x,\Delta y,\Delta z)$ on the $x$, $y$, and $z$ directions, and the changes of attitude angles $\mathbf{A}=(\Delta\phi,\Delta\theta,\Delta\psi)$ resulting from rotation on the three axes.
Specifically, $\mathbf{V}$ stands for linear velocity, which is the speed at which the UAV moves in real space and determines the location and speed of the UAV in the map. $\mathbf{A}$ represents the angular velocity, rotates around three axes ($x$, $y$ and $z$) and determines the attitude change and the turning speed of the UAV.
According to the Newton-Euler theorem, the nonlinear dynamics equations of a quadrotor UAV are [16],
$\left\{\begin{array}[]{l}\ddot{x}(t)=\frac{1}{m}U_{1}(t)(\sin\phi\sin\psi+\cos\phi\sin\theta\cos\psi)\par\\ \ddot{y}(t)=\frac{1}{m}U_{1}(t)(-\sin\phi\sin\psi+\cos\phi\sin\theta\sin\psi)\par\\ \ddot{z}(t)=\frac{1}{m}U_{1}(t)(\cos\phi\cos\theta)-g\\ \ddot{\phi}(t)=\frac{1}{I_{x}}\left(lU_{2}(t)-\dot{\theta}\dot{\psi}\left(I_{z}-I_{y}\right)\right)\par\\ \ddot{\theta}(t)=\frac{1}{I_{y}}\left(lU_{3}(t)-\dot{\phi}\dot{\psi}\left(I_{x}-I_{z}\right)\right)\par\\ \ddot{\psi}(t)=\frac{1}{I_{z}}\left(U_{4}(t)-\dot{\phi}\dot{\theta}\left(I_{y}-I_{x}\right)\right)\end{array}\right.$ (2)
where the left hand side of each equation represents the second order derivative of the relevant variable with respect to $t$. $I_{x}$, $I_{y}$, and $I_{z}$ are the rotational inertia of the UAV body around its own $x$, $y$, and $z$ axes, respectively. $I_{x}$ and $I_{y}$ are approximated based on the assumption of structural symmetry of the quadrotor UAV. Meanwhile, $l$ is the distance from the center of mass of the body to the center of the rotor, which is also equal to half of the airframe wheelbase.
Fig. 2 shows the schematic diagram of a quadrotor UAV in motion. Two pairs of motors $(M1,M4)$ and $(M2,M3)$ rotating in opposite directions are used to eliminate counter-torque. Moreover, $m$ is the mass of the UAV. $g$ is the acceleration of gravity, $I_{x}$, $I_{y}$, and $I_{z}$ are the rotational inertia of the UAV body around $x$, $y$, and $z$ axes, respectively.
Figure 2: Schematic diagram of a quadrotor UAV in motion
The input and output quantities interact with each other to change the UAV motion by adjusting the rotational speed of the four rotors (propellers). The trusts $F_{i}(t),(i=1,2,3,4)$ provided by the motors are directly proportional to the square of the angular velocity of the rotors (propellers) of the corresponding motors $\omega_{i}(t),(i=1,2,3,4)$, namely,
$F_{i}(t)=C_{T}(t)\omega_{i}(t)^{2},$ (3)
where $C_{T}(t)$ is the trust coefficient of the propellers. The torque representing the magnitude of the torque to overcome the air resistance is given by
$M_{p}(t)=C_{M}(t)\omega_{i}(t)^{2},$ (4)
with $C_{M}(t)$ as the torque coefficient.
For more details on principles related to attitude control and the calculation of the dynamics input $\mathbf{U}$, see Appendix -A.
III-C Rotor Propeller Design of Quadrotor UAV
Figure 3: Schematic of quadrotor UAV propeller hardware.
We now introduce the factors that would affect $C_{T}(t)$ and $C_{M}(t)$. The propeller is the power source of the UAV. Fig. 3 shows the structure of the quadrotor UAV propeller, where the propeller radius $R$ refers to the sum of the blade length $r$ and the radius of the mounting interface $r_{0}$. The blade mounting angle is the angle between the propeller blades and the plane of the UAV airframe, denoted by $\gamma$. Along with the blade width $P_{w}$, thickness $P_{t}$, we define the torque coefficient as,
[…]
where $N_{B}$ is the number of propeller blades, and $Q_{\gamma}(t)$ is determined by the propeller blade mounting angle and yaw angle in the current state. $\rho_{a}(t)$ is the air density determined by the atmospheric pressure at the location of the UAV and the absolute temperature, specifically,
$\rho_{a}(t)=\frac{P_{g}\cdot M_{g}}{R_{g}\cdot T(t)},$ (6)
where $P_{g}$ is the standard atmosphere, $M_{g}$ is the molecular weight of gas, $R_{g}$ is the gas constant, $T(t)$ is the outside ambient temperature and $z(t)$ is the height of flight at the location of the UAV. It is noteworthy that $M_{g}$ and $R_{g}$ are constants in the ideal gas state, and their values are given in Table V.
The values of propeller radius $R$ and airframe wheelbase $2l$ need to be set properly, in order to avoid collisions between neighboring propellers and underpowered situations. We illustrate the structural relationship in Fig. 4. In a symmetrically structured quadrotor configuration, the diagonal separation distance between neighboring propellers is $\sqrt{2}l$, and the total width of the two propellers ($2R$) must be less than this distance to avoid collision, namely $2R<\sqrt{2}l$. On the other hand, a too small $R$ (less than $l/3$) will not provide enough lift for the UAV according to (3). Thus, there is a constrained relationship between $R$ and $l$, namely $\frac{\sqrt{2}}{2}l\geqslant R\geqslant\frac{1}{3}l$.
Figure 4: Propeller radius in relationship to wheelbase for quadrotor UAV.
III-D Transmission Rate
We define a tuple $(m,n)$ to represent a transmission from $m$ to $n$, where $m$ and $n$ can be $j\in\{1,2,\cdots,K\}$ (representing the $j$th MD), $u$ (representing the UAV), or $D$ (representing the DC).
We consider $\zeta_{j}$ to be a parameter determined by the carrier frequency $f_{c}$ and obstacle density, corresponding to the proportion of the Line of Sight (LoS) path obstructed between the $j$th MD and the UAV, while the elevation angle of the link is $\theta_{j}$. The probability to establish a LoS connection between the $j$th MD and the UAV is,
$P^{\text{LoS}}_{j}(t)=\frac{1}{1+\zeta_{j}\cdot\exp\left(-\left(\theta_{j}-\zeta_{j}\right)\right)}.$ (7)
The Signal-to-Interference-plus-Noise Ratio (SINR) for $(m,n)$ at the $t$th timeslot is,
$\xi_{(m,n)}(t)=\frac{p_{j}(t)}{L_{(m,n)}(t)\cdot\left(I_{(m,n)}(t)+\sigma_{T}^{2}\right)},$ (8)
where $p_{j}(t)$ is the transmission power of the $j$th MD at the $t$th timeslot. $L_{(m,n)}(t)=P^{\text{LoS}}_{m}(t)\cdot\left(\frac{4\pi f_{c}}{c}d_{(m,n)}(t)\right)$ is the path loss, $c$ and $d_{(m,n)}(t)$ stand for the speed of light and the distance between $m$ and $n$, respectively. $I_{(m,n)}(t)$ refers to the interference caused by concurrent transmissions, and $\sigma_{T}^{2}$ is the thermal noise.
We can then obtain the effective transmission rate for $(m,n)$ at $t$th timeslot as,
$r_{(m,n)}(t)=W_{(m,n)}\log_{2}\left(1+\xi_{(m,n)}(t)\right),$ (9)
where $W_{(m,n)}$ is the bandwidth of the wireless link.
III-E Energy Consumption
We consider that the total energy consumption includes the consumption for the transmission and computation during the task offloading process, as well as those for the movement of the UAV. Hereafter, we use the binary variables $x_{ij}^{\text{MD}}(t)$, $x_{ij}^{\text{UAV}}(t)$, and $x_{ij}^{\text{DC}}(t)$ to specify the assignment of the $i$th task from the $j$th MD at the $t$th timeslot. A value of $1$ indicates that the task is assigned to the corresponding location.
III-E1 UAV movement
We consider that the motion state of the UAV is directly related to the rotational speed of the four DC motors. Thus, the energy consumption due to UAV movement at the $t$th timeslot is,
$E^{\mathrm{MOV}}(t)=\sum_{i=1}^{4}\left(I_{i}(t)\cdot L\right)=\sum_{i=1}^{4}\left(\frac{V_{m}-\omega_{i}(t)P_{m}}{R_{m}}\cdot L\right),$ (10)
where $I_{i}(t)$ is the motor current, which can be calculated by (28) and (2) given that $U$ and $\omega_{i}(t)(i=1,2,3,4)$ (the linear speed of $i$th motor rotation (r/min) at the $t$th timeslot) are obtained by our proposed attitude control method. $V_{m}$ is the motor rated voltage, $R_{m}$ is the rated resistance of the UAV motor circuit, and $P_{m}$ is the electric potential constant determined by the motor structure.
III-E2 Energy consumption for transmission
If a task is assigned to the UAV, two segments of transmission will be incurred, namely from the MD to the UAV and the UAV back to the MD. On the other hand, if a task is assigned to the DC, two additional segments, namely from the UAV to the DC and from the DC back to the UAV, will be incurred. Therefore, energy consumption at the $t$th timeslot for transmitting the $i$th task from the $j$th MD is,
$e_{ij}^{\text{TR}}(t)=$ $\left(x_{ij}^{\text{UAV}}(t)+x_{ij}^{\text{DC}}(t)\right)\left(\frac{p_{j}(t)s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right)$ (11)
$+x_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{p_{D}(t)s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right),$
where $C_{j}(t)$ is the number of channels assigned to the $j$th MD.
III-E3 Energy consumption for computation
We consider that the additional energy consumption for computing a finite number of tasks in the DC is negligible as the DC is assumed to be always active in processing tasks from different sources. Therefore, the energy consumption for computing $i$th task of the $j$th MD can be obtained by
$e_{ij}^{\mathrm{COMP}}(t)=x_{ij}^{\text{UAV}}(t)\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}+x_{ij}^{\text{MD}}(t)\delta_{j}s_{ij}(t)c_{ij}\left(f_{j}(t)\right)^{2},$ (12)
where $\delta_{u}$ and $\delta_{j}$ are the effective computing capacities of the UAV, and the $j$th MD, respectively.
III-E4 Overall Energy Consumption
based on the discussions above, the overall energy consumption at the $t$th timeslot can be expressed as,
$\mathbb{E}(t)=\sum_{j=1}^{K}\sum_{i=1}^{N_{j}}\left[e^{\mathrm{TR}}_{ij}(t)+e^{\mathrm{COM}}_{ij}(t)\right]+E^{\mathrm{MOV}}(t),$ (13)
where $N_{j}$ is the total number of tasks for the $j$th UD.
III-F Delay
We consider three delay components: transmission, computation, and queuing.
III-F1 Transmission delay
the transmission delay for the $i$th task from the $j$th MD at the $t$th timeslot is,
$d_{ij}^{\text{TR}}(t)$ $=x_{ij}^{\text{DC}}(t)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right)$ (14)
$+\left(x_{ij}^{\text{UAV}}(t)+x_{ij}^{\text{DC}}(t)\right)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right).$
III-F2 Computation delay
the delay for computing the $i$th task from the $j$th MD can be expressed as,
$d_{ij}^{\mathrm{COMP}}(t)=x_{ij}^{\text{UAV}}(t)\frac{s_{ij}(t)c_{ij}}{f(t)}+x_{ij}^{\text{MD}}(t)\frac{s_{ij}(t)c_{ij}}{f_{j}(t)}.$ (15)
Note that as the computing resources at the DC are sufficient for all practical purposes, we assume that the computation delay at the DC is negligible.
III-F3 Queuing delay
Consider that the arrival rate for transmission from the $j$th MD to UAV is $\lambda_{j}$, and the processing rate at the UAV for receiving transmissions is $[L\cdot C_{j}(t)\cdot r_{j,u}(t)]/s_{j}(t)$, where $s_{j}(t)=\sum_{i=1}^{N_{j}}s_{ij}(t)$ is the sum of tasks data size from that MD. If the total arrival rate in larger than the processing rate, a queuing delay for transmission will be incurred [13], that is,
$d_{ij}^{\mathrm{Q}}(t)=\frac{\lambda_{j}\left(1-p_{j}\right)s_{j}(t)^{2}}{L\cdot C_{j}(t)r_{j,u}(t)\cdot\left(L\cdot C_{j}(t)r_{j,u}(t)-\lambda_{j}\left(1-p_{j}\right)s_{j}(t)\right)}.$ (16)
Similarly, if the total sizes of arriving tasks exceed the computation capacity of the component, queuing delay for computation would be incurred for a proportion of the tasks. The expected queuing delay of the $i$th task from the $j$th MD executed at UAV at $t$th timeslot is,
$d_{ij}^{\mathrm{UAVQ}}(t)=\frac{\bar{Q}}{\sum_{j=1}^{K}\lambda_{j}\left(1-p_{j}\right)}-\frac{\sum_{j=1}^{K}\lambda_{j}\left(1-p_{j}\right)s_{j}(t)c_{j}}{\tau f(t)\sum_{j=1}^{K}\lambda_{j}\left(1-p_{j}\right)},$ (17)
where $\bar{Q}$ is a parameter related to the computing capacity of the UAV.
III-F4 Total delay
The total delay is obtained by summing all delay components mentioned earlier, that is,
$\mathbb{D}(t)=\sum_{j=1}^{K}\sum_{i=1}^{N_{j}}\left(d_{ij}^{\mathrm{TR}}(t)+d_{ij}^{\mathrm{COMP}}(t)+d_{ij}^{\mathrm{UAVQ}}(t)+d_{ij}^{\mathrm{Q}}(t)\right).$ (18)
IV Joint Optimization Problem
IV-A Attitude Control
IV-A1 Classical PID Control
The classical PID control algorithm is a widely used algorithm for UAV attitude control [17]. It considers the following three control parts:
  • Proportional control (P): the error of the controller input is amplified by a certain proportion. This constitutes the basic control part;
  • Integral control (I): an integral term is introduced for correction of the proportional control to maintain the system stability. Its coefficient will approach 00 as time goes by;
  • Differential control (D): a differential term based on the rate of change of the input error is added. The purpose is to improve the response accuracy by predicting the future trend of the input error.
The classical PID control is described by the following equation,
[…]
where $k_{p}$, $k_{i}$ and $k_{d}$ are the proportional, integral and differential coefficients, respectively. The output of the PID control system, $U(t)$, will affect $\omega,\theta,\phi,\psi$, which will in turns have an impact on the power consumption of the UAV (see Section III-B and Appendix -A for details).
IV-A2 Fuzzy PID Control
A major shortcoming of the classical PID control is that, as the coefficients $k_{p}$, $k_{i}$, and $k_{d}$ are constants, it cannot adapt to dynamic environments and inter-dependent parameters as in UAV assisted fog computing scenarios. Therefore, we propose the fuzzy PID control algorithm, a nonlinear control technique with adaptive capability based on the Fuzzy set theory [18].
The fuzzy control process is demonstrated in Fig. 5. The key part of the process is the fuzzy controller at the top of Fig. 5), which determines $\Delta k_{p}$, $\Delta k_{i}$, and $\Delta k_{d}$, the adjustments in control coefficients, based on the yaw angle error $E$ and the error rate $EC$ from the previous estimation. The specific rules are shown in Table III. Based on the demand for Mamdani reasoning method in solving fuzzy PID problems [40], we establish the the allowable value of $E$, $EC$, $\Delta k_{p}$, $\Delta k_{i}$, and $\Delta k_{d}$ in Table IV.
TABLE III: FUZZY PID CONTROL RULES
NB: Negative Big ($-3$)  NM: Negative Middle ($-2$)   NS: Negative Small ($-1$)  ZO (0): zero
PB: positive big ($3$)   PM: positive middle ($2$)   PS: positive small ($1$)
  $EC$
$E$ NB NM NS ZO PS PM PB
[…]
NB PB/NB/PS PB/NB/NS PM/NM/NB PM/NM/NB PS/NS/NB ZO/ZO/NM ZO/ZO/PS
NM PB/NB/PS PB/NB/NS PM/NM/NB PS/NS/NM PS/NS/NM ZO/ZO/NS NS/PS/ZO
NS PM/NB/ZO PM/NM/NS PM/NS/NM PS/NS/NM ZO/ZO/NS NS/PM/NS NS/PM/ZO
ZO PM/NM/ZO PM/NM/NS PS/NS/NS ZO/ZO/NS NS/PS/NS NM/PM/NS NM/PM/ZO
PS PS/NM/ZO PS/NS/ZO ZO/ZO/ZO NS/PS/ZO NS/PS/ZO NM/PM/ZO NM/PB/ZO
PM PS/ZO/PB ZO/ZO/PM NS/PS/PM NM/PS/PM NM/PM/PS NM/PB/PS NB/PB/PB
PB ZO/ZO/PB ZO/ZO/PM NM/PS/PM NM/PM/PM NM/PM/PS NB/PB/PS NB/PB/PB
TABLE IV: FUZZY PID PARAMETER SETTINGS
Parameters Values Parameters Values Parameters Values
$E$, $EC$6, -5, -4, -3, -2, -3, -2, -1, -3, -2, -1,
-1, 0, 1, 2, $\Delta k_{p}$, $\Delta k_{d}$ 0, 1, $\Delta k_{i}$ 0, 1,
3, 4, 5, 6 2, 3 2, 3
$\left\{\begin{array}[]{l}k_{p-Fuzzy}=k_{p}+\Delta k_{p}\\ k_{i-Fuzzy}=k_{i}+\Delta k_{i}\\ k_{d-Fuzzy}=k_{d}+\Delta k_{d}.\end{array}\right.$ (20)
Figure 5: Schematic structure of Fuzzy PID control system for quadrotor UAV.
IV-B Joint optimization model
We consider that the operational efficiency cost of the network is determined by the weighted sum of the energy and delay components. For the $t$th timeslot, the objective function is defined for all tasks originating from the MDs, taking into account both total energy consumption defined in (13) and total time required to complete all tasks defined in (18). The joint optimization problem can then be defined as,
$\begin{array}[]{ll}\underset{x_{ij}}{\operatorname{Minimize}}&\mathbb{S}(t)=\mathbb{D}(t)+\epsilon\cdot\mathbb{E}(t)\\ \text{S.t.}&\left(\complement_{1}\right)\hskip 5.69046pt0\leq v(t)\leq v^{\max}\\ &\left(\complement_{2}\right)\hskip 5.69046pt0\leq z(t)\leq z^{\max}\\ &\left(\complement_{3}\right)\hskip 5.69046pt\sum_{j=1}^{K}C_{j}(t)\leq N_{c}\\ &\left(\complement_{4}\right)\hskip 5.69046ptf^{\min}\leq f(t)\leq f^{\max}\\ &\left(\complement_{5}\right)\hskip 5.69046ptf_{j}^{\min}\leq f_{j}(t)\leq f_{j}^{\max}\\ &\left(\complement_{6}\right)\hskip 5.69046ptp^{\min}\leq p(t)\leq p^{\max}\\ &\left(\complement_{7}\right)\hskip 5.69046ptp_{j}^{\min}\leq p_{j}(t)\leq p_{j}^{\max}\\ &\left(\complement_{8}\right)\hskip 5.69046ptx_{ij}^{\text{MD}}(t)+x_{ij}^{\text{UAV}}(t)+x_{ij}^{\text{DC}}(t)=1\\ &\left(\complement_{9}\right)\hskip 5.69046ptx_{ij}^{\text{MD}}(t),x_{ij}^{\text{UAV}}(t),x_{ij}^{\text{DC}}(t)\in\{0,1\}\\ &\left(\complement_{10}\right)\sum_{t=0}^{T}\left(E^{\text{UAV}}(t)\right)\leq B_{c}\\ &i\in\{1,\ldots,N_{j}\},j\in\{1,\ldots,K\},\end{array}$ (21)
The parameter $\epsilon$ is a weighting factor in the objective function. One can tune this weighting factor to adjust the relative importance of delay and energy consumption, flexibly catering to different magnitudes and ranges of the two measurements, as well as different requirements under various scenarios in UAV-assisted fog computing and reflecting the preference in terms of the trade-off between the two metrics. In terms of the constraints, $\complement_{1}$ to $\complement_{3}$ were described in detail previously when they first appeared. $\complement_{4}$ to $\complement_{7}$ ensure that the processing frequency and transmission power of UAV and MDs are always within the effective range. $\complement_{8}$ and $\complement_{9}$ guarantee that all tasks are at exactly one place during a timeslot. $\complement_{10}$ restricts that the total energy consumption of the UAV during the whole period of $T$ does not exceed the capacity of its battery, where $E^{\text{UAV}}(t)$ denotes the energy consumption of the UAV at the $t$th timeslot,
$E^{\text{UAV}}(t)$ $=E^{\text{MOV}}(t)+\sum_{j=1}^{K}\sum_{i=1}^{N_{j}}\left[x_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}\right)\right.$ (22)
$\left.+x_{ij}^{\text{UAV}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}+\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}\right)\right].$
Note that the processing frequencies $f(t)$, $f_{j}(t)$, and transmission powers $p(t)$, $p_{j}(t)$ are dependent on the decision variables $x_{ij}(t)$. Numerous scholarly works provide comprehensive analyses of these intricate interactions, e.g., [13].
V Algorithms
The problem (21) possesses an intrinsic nonconvex nature. We divide the problem into two subproblems, one focusing on generating an initial trajectory and altitude, and the other on task assignment and resource allocation, as well as further tuning the trajectory. We propose two algorithms to address the two subproblems, respectively, with enhanced decoupling and safety values mechanism incorporated in the second algorithm to overcome the deadlock issue in conventional heuristic algorithms.
An overview of the steps in our proposed framework to address the joint optimization problem is shown in Fig. 6. We will apply an Ant Colony System based algorithm (Algorithm 2, to be described in detail later) to determine an initial trajectory based on the terrain of the area and the position of MDs. Then, the attitude control mechanism described in Section IV-A will be invoked to determine the as well as tune the trajectory. Finally, a Particle Swarm Optimization based algorithm (Algorithm 1, which was initially proposed in the conference version [14]) will determine the optimal task assignments and resource allocations. We will also analyze the convergence performance of the proposed algorithms at the end of the section.
Figure 6: A flowchart demonstration of the components in our proposed framework.
V-A PSO for Task Assignment and Resource Allocation
We first introduce an algorithm based on the idea of Particle Swarm Optimization (PSO), for assigning tasks and allocating resources given that the trajectory is already known. The steps of our proposed algorithm are demonstrated in Algorithm 1, which we refer the readers to the conference version of this paper [14] for more detailed explanations. We categorize the job assignment decision, transmission power, and processing frequency as a collective entity $s_{h}=(x_{ij},p_{j}(t),f_{j}(t))$ within the particle swarm. Here, $h\in\{1,...,H\}$. Next, the determination of the particle group’s position is carried out by uniformly sampling a set of particle groups denoted by $H\in\mathbb{N_{+}}$. In order to search for the optimal solution, the velocity of the particle groups is initialized and subsequently updated using the method in [41]. The algorithm terminates when the discrepancy between the outcomes of $20$ successive rounds is less than a specified threshold $\eth$, and then outputs the minimal value $\mathbb{S}^{*}$, the optimal decision for job assignment $x_{ij}^{*}$, the transmission power $p_{j}^{*}(t)$, and the processing frequency $f_{j}^{*}(t)$.
Algorithm 1 Task Assignment, Power and Frequency Allocations based on Particle Swarm Optimization
1:The length of an interval $t$, $p^{\mathrm{max}}_{j}(t)$; $p^{\mathrm{min}}_{j}(t)$; $f^{\mathrm{max}}_{j}(t)$; $f^{\mathrm{min}}_{j}(t)$; $F_{A1}$; $F_{A2}$; $F_{I}$
2:$\mathbb{S}^{*}(t)$; $x_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t)$
3:for $h=1$ to $H$ do
4:     Initialize particles’ position $G_{h}$
5:     Initialize particles’ velocity $V_{h}$
6:     Initialize $pBest_{h}\longleftarrow G_{h}$
[…]
8:for $h=1$ to $H$ do
9:     Update $V_{h}$ and $G_{h}$ by acceleration factors $F_{A1}$, $F_{A2}$ and $F_{I}$
10:     if $fit(G_{h})<fit(pBest_{h})$ then
11:         $pBest_{h}\longleftarrow G_{h}$
12:         if $fit(pBest_{h})<fit(gBest)$ then
13:              $gBest(k)\longleftarrow pBest_{h}$
14:     $k\longleftarrow(k+1)$
15:while $\lvert gBest(k+1)-gBest(k)\rvert<\eth$ do
16:     $\mathbb{S}^{*}(t)\longleftarrow fit(gBest)$
17:     $(x_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$ $\longleftarrow gBest$
18:Output $\mathbb{S}^{*}(t)$; $(x_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$
V-B ACS-DS for Trajectory Planning: Motivation and Concept
We now present the concept and outlines of the ACS-DS (ACS with Decoupling and Safety values) algorithm to generate the optimal UAV trajectory to minimize $E^{\mathrm{MOV}}$. Then, we will prove that ACS-DS always converges to the optimal solution in polynomial time in subsequent subsections.
We first divide the 3D space into discrete grids, with the center of each unit of the grid serving as a waypoint, that is, the position that the UAV will pass. The 3D terrain is generated randomly, and we denote the area at and below terrain surfaces, referred as the no-fly zone, by $\boldsymbol{O}=\{(x^{o}_{n},y^{o}_{n},z^{o}_{n})\}$, where $n\in\{1,2,...,N_{o}\}$, and $N_{o}$ is the total number of grids in the no-fly zone. Note that the classical ACS is not applicable to our three-dimensional trajectory planning problem as it struggles to converge in path searching in large three-dimensional spaces, and is likely to be trapped in local optima.
To overcome these issues and obtain the optimal trajectory with minimum cost in an efficient manner, we propose the ACS-DS, by incorporate the two special mechanisms to the classical ACS. We now elaborate these two mechanisms.
V-B1 Safety Values
We enhance the heuristic function of the ACS-DS by adding a security value at each step based on the number of currently known the numbers of feasible and infeasible waypoints. Specifically, we determine the safety value of a waypoint based on the proportion of known feasible waypoints in the next available position for that waypoint. The safety value from waypoint $\mu$ to $\nu$ is calculated as $\kappa_{\mu\nu}=\frac{N_{f}-N_{\mu\nu}}{N_{f}}$, where $N_{f}$ denotes the total number of waypoints in a preset constant range $\left[0,R_{f}\right]$, and $N_{\mu\nu}$ denotes the number of infeasible waypoints in the range of $N_{f}$ over direction from $\mu$ to $\nu$. Safety values are updated using a similar rule to pheromones, and when choosing its next move, the ant would add the safety values to the pheromone levels for all possible actions. We will show in the results section that this method can reduce the running time and improve the operation efficiency compare to the classical ACS algorithm.
V-B2 Decoupling
In ACS, when an ant reaches a point with no further viable options, it becomes trapped in a deadlock. To address this issue, we introduce a mechanism that allows the ant to escape the deadlock. Specifically, when any of following rule is satisfied during the ant’s movement, the Decoupling mechanism will be triggered to perform a backtracking behavior with depth (step size) $D_{b}$ to find alternative directions:
  • An ant repeats a closed cycle consisting of two or more waypoints over multiple consecutive timeslots.
  • A sudden drop in the amount of pheromone and safety values occurs, as indicated by the fact that the $\kappa$ of the currently selected waypoint is less than half that of the previous waypoint.
  • An ant falls into a local optimum and undetected in the early stage, that is, the ant has not triggered backtracking behavior for more than 25 consecutive waypoints in the first one-third of all iterations.
It is worth noting that the third rule is only considered in the early stages of the algorithm (first one-third of all iterations) to avoid excessive backtracking that reduces the convergence speed of the algorithm. This mechanism, by dynamically adjusting the pheromone levels and allowing the ant to backtrack and explore new possibilities, ensures that the ACS-DS remains robust and capable of finding optimal trajectories even in complex and challenging scenarios.
V-C ACS-DS: Detailed Steps
We let $m$ be the total number of ants in the colony, and $\sigma=V_{P}+\kappa$ be the guidance factor. We further denote $\sigma_{\mu\nu}(h)$ as the sum of pheromone values and safety values between the neighboring waypoints $\mu$ and $\nu$ in the $h$th iteration. The initial pheromones on each edge are equal, namely $\sigma_{\mu\nu}(0)=C$ for all $\mu$ and $\nu$. We denote by $s^{*}$ the global best path obtained by the algorithm, corresponding to the minimal-cost trajectory in our UAV optimization problem. Next, for each ant $k\in\{1,2\ldots,m\}$ in the colony, we initialize the pheromone $V_{P0}$, the safety value $\kappa_{0}$, and the heuristic value $\eta$. We also define the evaporation rate $\rho\in(0,1)$ representing the degree to which the guidance factor $\sigma_{\mu\nu}(h)$ decays with iterations. Finally, we define $S_{\iota}$ as the set of waypoints that ant $\iota$ can pass next, and $p_{\mu\nu}^{\iota}(h)$ as the probability that ant $\iota$ moves from position $\mu$ to $\nu$ in the $h$th iteration. The exact value of $p_{\mu\nu}^{\iota}(h)$ is jointly determined by the guidance factor and heuristic value at the waypoint, as in the following equation,
$p_{\mu\nu}^{\iota}(h)=\begin{cases}\frac{\sigma_{\mu\nu}^{\alpha}(h)\eta_{\mu\nu}{}^{\beta}(h)}{\sum_{r\in S_{\iota}}\sigma_{ir}^{\alpha}\eta_{ir}{}^{\beta}(h)}&\nu\in S_{\iota},\\ \quad\quad\quad 0&\text{ otherwise.}\end{cases}$ (23)
Detailed steps of ACS-DS are presented in Algorithm 2.
Algorithm 2 ACS-DS: ACS-based Trajectory Planning with Decoupling and Safety Values Mechanisms
1:Positions of all MDs $(x_{j},y_{j},z_{j})$; $\rho$; $\eta$; $V_{P0}$; $\kappa_{0}$; terrain surface with no-fly zone $(x^{o}_{n},y^{o}_{n},z^{o}_{n})\in\boldsymbol{O}$
2:$E^{\mathrm{MOV}*}(t)$; Trajectory of the UAV $(x_{\iota}(t),y_{\iota}(t),z_{\iota}(t))$
3:for $h=1$ to $H$ do
4:     Randomly set the initial coordinate of the UAV, $(x(0),y(0),z(0))$
5:     while $(x(0),y(0),z(0))\notin\boldsymbol{O}$ do
6:         for each edge do
7:              Set initial pheromone, and calculate the initial safety values
8:         for each ant $\iota$ do
9:              $(x_{\iota}(h),y_{\iota}(h),z_{\iota}(h))$ = $(x(0),y(0),z(0))$
10:              for each edge $(\mu,\nu)$ do
11:                  Choose the next coordinate with probability $p_{\mu\nu}^{\iota}(h)$ by $\sigma$ and $\eta$
12:                  while $(x_{\iota}(h+1),y_{\iota}(h+1),z_{\iota}(h+1))\notin\boldsymbol{O}$ do
13:                       Output $(x_{\iota}(h+1),y_{\iota}(h+1),z_{\iota}(h+1))$
14:              Compute and output the length $\sum_{t=1}^{T}d_{u}(t)$ of the path by the $\iota$th ant and $E_{u}$
15:              for each edge $(\mu,\nu)$ do
16:                  Update $V_{P_{\mu\nu}}$ and $\kappa_{\mu\nu}$ by $\rho$
17:                  Update $\sigma_{\mu\nu}(h)$ by $V_{P_{\mu\nu}}$ and $\kappa_{\mu\nu}$
18:Compute and output $E^{\mathrm{MOV}*}(t)$ by (10).
V-D ACS-DS: Proof of convergence
We now prove that Algorithm 2 will eventually converge, starting by the following proposition.
Proposition 1.
In Algorithm 2, for the guidance factor $\sigma_{\mu\nu}$ on any edge $(\mu,\nu)$ generated by the ants during the searching process, there exists a maximum value $g(s^{*})$ as $h\to\infty$.
Proof.
For ACS-DS, the local update of the guidance factor for edge $(\mu,\nu)$ after the completion of each round of iteration can be represented as $\sigma_{\mu\nu}(h+1)=(1-\varphi)\cdot\sigma_{\mu\nu}(h)+\varphi\cdot\Delta\sigma_{\mu\nu}(h)$, while the global guidance factor is updated in a similar way, that is, $\sigma(h+1)=(1-\varphi)\cdot\sigma(h)+\varphi\cdot\Delta\sigma(h)$. Here, $\Delta\sigma_{\mu\nu}(h)=\sigma_{\mu\nu}(h)-\sigma_{\mu\nu}(h-1)$ is the increment of guidance factor on edge $(\mu,\nu)$ at the $h$th iteration.
$\sigma_{\mu\nu}(1)=(1-\varphi)\cdot\sigma_{\mu\nu}(0)+\varphi\cdot\Delta\sigma_{\mu\nu}(0)$ (24)
$\cdots$
$\sigma_{\mu\nu}(h)=(1-\varphi)^{h}\cdot\sigma_{\mu\nu}(0)+(1-\varphi)^{h-1}\cdot\varphi\cdot\Delta\sigma_{\mu\nu}(1)+\cdots$
$\quad\quad\quad\quad+(1-\varphi)\cdot\varphi\cdot\Delta\sigma_{\mu\nu}(h-1)+\varphi\cdot\Delta\sigma_{\mu\nu}(h)$
$\hskip 24.75375pt=\sum_{q=1}^{h}\left((1-\varphi)^{h-q}\cdot\varphi\cdot\Delta\sigma_{\mu\nu}(q)+(1-\varphi)^{h}\cdot\sigma_{\mu\nu}(0)\right)$
$\lim_{h\rightarrow\infty}\sigma_{\mu\nu}^{\max}(h)=\lim_{h\rightarrow\infty}\left(\sum_{q=1}^{h}\left((1-\varphi)^{h-q}\cdot\varphi\cdot\Delta\sigma_{\mu\nu}(q)+\cdots\right)\right)$
$\quad\quad\quad\quad\quad=g(s^{*})<\infty.$
Thus, the guidance factor on each edge is bounded from the above by $g(s^{*})$. ∎
After the first optimal solution is found, the guidance factor on the elements belonging to an optimal solution is guaranteed to be no less than that on other elements as we have a sufficient number of subsequent generations. That is, the guidance factor on any element not belonging to an optimal solution will keep decreasing until it is no larger than the guidance factor on the elements belonging to an optimal solution. In mathematical terms, we have the following corollary.
Corollary 1.
As $h\to\infty$, it is always true that $\sigma_{\mu\nu}(h)\geq\sigma_{\mu^{\prime}\nu^{\prime}}(h)$, if $(\mu,\nu)\in s^{*}$ and $(\mu^{\prime},\nu^{\prime})\notin s^{*}$.
Proof.
Suppose it takes $h^{*}$ generations to find the first optimal solution. We assume that for a certain $(\mu,\nu)\in s^{*}$, there exists a $(\mu^{\prime},\nu^{\prime})\notin s^{*}$, $\sigma_{\mu\nu}\left(h^{*}\right)<\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}\right)$. Then, based on the guidance factor update rule in ACS-DS, by the $h^{*}+h^{\prime}$ generation, the guidance factor on $(\mu^{\prime},\nu^{\prime})$ can be derived as
$\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}+h^{\prime}\right)=\max\left\{\sigma_{\mu\nu}\left(h^{*}\right),(1-\rho)^{h^{\prime}}\cdot\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}\right)\right\}$ (25)
Therefore,
$\lim_{h\rightarrow\infty}\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}+h^{\prime}\right)$ $=\max\lim_{h\rightarrow\infty}\left\{\sigma_{\mu\nu}\left(h^{*}\right),\left[(1-\rho)^{h^{\prime}}\cdot\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}\right)\right]\right\}$ (26)
$=\max\left\{\sigma_{\mu\nu}\left(h^{*}\right),0\right\}$
$\leq\sigma_{\mu\nu}\left(h^{*}\right)$
Proposition 1 and Corollary 1 jointly prove that, after a sufficient number of iterations, the guidance factor on an optimal path is bounded and no less than those on other paths. Noticeably, the backtracking behavior caused by the decoupling mechanism in ACS-DS will only increase the number of iterations for the current waypoint, but does not affect the updating of the guidance factor and the convergence to the optimal solution as in the original ACS. Therefore, the ACS-DS algorithm always converges to the global optimal solution.
V-E ACS-DS: Complexity analysis
We now analyze the complexity of Algorithm 2 when a total of $n$ users initiate computation requests in the current timeslot. Assume that in the worst case, Algorithm 2 triggers backtracking at every waypoint for each ant upfront, the path construction time complexity for $m$ ants is $O\left(m\cdot n^{2}\right)$. The complexity of each update of the guidance factor, and each calculation of the transfer probability is $O\left(n^{2}\right)$. If the algorithm is finished after $H$ iterations, the complexity of the entire algorithm is $O\left(H\cdot m\cdot n^{2}+H\cdot n^{2}\right)$. Since the magnitude of $m$ is similar to $n$, the complexity of the ACS-DS algorithm in completing all the tasks of the network is approximately $O\left(H\cdot n^{3}\right)$, which means that Algorithm 2 has a polynomial time complexity.
VI Performance Evaluation
In this section, we present numerical results of simulations on systems with various settings to assess the efficacy and versatility of our proposed solutions.
VI-A Experiment setup
We consider a three-dimensional spatial domain with dimensions of $S\times S\times Z$. $K$ MDs are uniformly and randomly distributed in the space. Values of key system parameters are listed in Table V.
TABLE V: Values of relevant system parameters in the experiments
Parameter Value Parameter Value
$S$ 50000 $\mathrm{~{}m}$ $Z$ 3000 $\mathrm{~{}m}$
[…]
$m$ 80 $\mathrm{~{}kg}$ $a$ 5.28 $\mathrm{~{}m}/\mathrm{s}^{2}$
$g$ 9.81 $\mathrm{~{}m}/\mathrm{s}^{2}$ $l$ 1150 $\mathrm{~{}mm}$
$K$ 50 $R$ 512 $\mathrm{~{}mm}$
$N_{c}$ 40 $P_{t}$ 7.5 $\mathrm{~{}mm}$
$N_{B}$ 2 $P_{w}$ 250 $\mathrm{~{}mm}$
$B_{c}$ 5000 $\mathrm{~{}Wh}$ $I_{x}$ 0.095
[…]
$z^{\max}$ 2500 $\mathrm{~{}m}$ $I_{z}$ 0.187
$C_{T}$ 1.483 ${}\left(\mathrm{N}\cdot\mathrm{m}\cdot\mathrm{min}^{2}\right)/\mathrm{r}^{2}$ $C_{M}$ 2.925 ${}\left(\mathrm{N}\cdot\mathrm{m}\cdot\mathrm{min}^{2}\right)/\mathrm{r}^{2}$
$M_{g}$ 29 $\mathrm{~{}g/mol}$ $R_{g}$ 8.314 $\mathrm{~{}J/(mol*K)}$
The values of parameters used in the algorithms described earlier in this paper are listed in Table VI. All experimental results presented in this section are based on the average of the 50 independent runs with the weighting factor $\epsilon$ randomly generated randomly within its domain for each run.
TABLE VI: Values of relevant algorithmic parameters
Parameter Value Parameter Value Parameter Value
$\rho$ 0.25 $V_{P0}$ 3.8 $\eta$ 2.5
$F_{A1}$, $F_{A2}$ 2.0 $F_{I}$ 0.65 $R_{f},D_{b}$ 200 ${}\mathrm{m}$
For the task assignment and resource allocation part, we adopt PSO (Algorithm 1), which has been demonstrated effective and robust in our earlier conference paper [14]), for all experiments in this section. For the attitude control and trajectory planning modules, we focus on evaluating the performance of our proposed ACS-DS (Algorithm 2) and ATC (Section IV-A). Specifically, we examine the following eight implementations and compare their performances in terms of overall operational efficiency cost.
  • RAN: All tasks are randomly assigned to the MD, the UAV, or the data center. The channels, processing frequencies and power allocations are generated randomly in their respective domains. No special mechanisms are applied for attitude control or trajectory planning.
  • PSO: The PSO algorithm (Algorithm 1) is used to determine task assignments, the processing frequencies, and transmission powers of MDs and the UAV. Allocation of channels in accordance with the methodology described in Section III. No special mechanisms are applied for attitude control or trajectory planning.
  • ACS: On top of PSO, trajectory of the UAV is planned by the classical ACS.
  • ACS $+$ ATC: On top of ACS, the mechanism described in Section IV-A is added to control the attitude of the UAV.
  • ACS-D: On top of PSO, trajectory of the UAV is planned by ACS with the decoupling mechanism.
  • ACS-D $+$ ATC: The mechanism described in Section IV-A is added to control the attitude of the UAV on top of ACS-D.
  • ACS-DS: On top of PSO. Trajectory of the UAV is planned by ACS with decoupling and safety values mechanisms in Algorithm 2.
  • ACS-DS $+$ ATC: The mechanism described in Section IV-A is added to control the attitude of the UAV on top of ACS-DS.
VI-B Numerical results
VI-B1 Optimal Propeller Parameters
Fig. 7 shows the relationship between the overall energy consumption and propeller blade parameters including the number of blades, and the radius, width, mounting angle, and thickness of each blade, given that all other parameters are fixed. From the results, we can infer that, for a symmetric quadrotor UAV with a total mass of $80~{}\mathrm{kg}$, four rotors and a maximum acceleration up to $5.28~{}\mathrm{m}/\mathrm{s}^{2}$, it is optimal to equip $2$ propeller blades with a radius of $500~{}\mathrm{mm}$, a width of $250~{}\mathrm{mm}$ and a thickness of $7.5~{}\mathrm{mm}$. The average energy consumption under configurations with optimal values of parameters can be reduced by more than $30\%$, compared to the average amount of 50 sets of random parameters generated uniformly within the respective allowable range of each parameter.
Figure 7: Relationship between overall energy consumption and propeller blades parameters.
VI-B2 Convergence Speed of Algorithms
Fig. 8 demonstrates the average convergence speeds of $200$ independent runs by the classical ACS, ACS-D and ACS-DS over $80$ iterations. Both mechanisms, decoupling and safety values, reduce the energy consumption of UAV movement to an extent. In addition, with classical ACS as the baseline, ACS-D improves the convergence speed to obtain the optimal trajectory by more than $7\%$, while the improvement by ACS-DS is more than $15\%$.
Figure 8: Convergence performances of trajectory planning algorithms.
VI-B3 Trajectory planning
Figure 9: A demonstration of trajectories planned by different mechanisms on a 3D spatial domain with terrain height generated randomly as a continuous surface.
We now generate the terrain height randomly as a continuous surface (no-fly zone) within the three-dimensional spatial domain as shown in Figs. 9 and 10, and consider that the MDs are placed directly on the ground conforming to the terrain height, in order to test the obstacle avoidance capability of our proposed trajectory planning and attitude control algorithms. We observe from the trajectories of the UAV in Figs. 9 and 10 that, both the additional mechanisms of decoupling and safety values in ACS and attitude control (ATC) have positive impacts on the final path. Specifically, ATC achieves more stable and precise flight control by optimizing the UAV rotor speed in real time based on the feedback of error and environmental information from the sensors. By jointly applying the ACS with the decoupling, safety values and attitude control, ACS-DS+ATC obtains the smoothest and shortest trajectory among all.
Figure 10: Demonstration of trajectories planned by different mechanisms on a 3D spatial domain with terrain height generated randomly as a continuous surface (TOP).
VI-B4 Operational Efficiency Cost
Finally, we present the results of the operational efficiency cost (combining delay and energy consumption) achieved by the eight implementations in Fig. 11. Again, we can observe that the decoupling mechanism and safety values in ACS-DS can significantly reduce the overall consumption compared to the classical ACS.
In addition, our proposed attitude control mechanism (ATC) can further improve the performance, irrespective of whether ACS, ACS-D, or ACS-DS is employed. Overall, ACS-DS+ATC can reduce the operational efficiency cost by more than $58\%$ and $35\%$ compared to RAN and classical ACS, respectively.
Figure 11: Comparison of operational efficiency cost by different methods.
VII Conclusions
In this paper, we have proposed a joint optimization framework to reduce the operational efficiency cost in UAV-assisted fog computing systems. Our proposed framework involves multiple modules, including quadrotor UAV attitude control, trajectory planning in a three-dimensional spatial domain with continuously varying terrain heights, and energy-efficient assignment of computing tasks to different components in the network. We have designed appropriate mechanisms or algorithms for each module in the framework, and integrated them together to obtain a holistic solution to improve the overall efficiency in UAV-assisted fog computing. Specifically, we have proposed a novel fuzzy PID control mechanism for effective attitude control, designed the ACS-DS algorithm that overcomes the convergence issue in conventional approaches for trajectory planning in three-dimensional domains, and a modified PSO algorithm to determine the optimal task assignment. Numerical results from a wide range of experiments have shown that our proposed framework can reduce the operational efficiency cost by over $35\%$ compared to existing approaches.
-A The Dynamics of a Quadrotor UAV
We consider two key components in the structure of a quadrotor UAV, namely the body frame and the earth frame. The body frame is commonly used for attitude control of the UAV, where the positive direction is the direction of ascent corresponding to the centers of the four motors.
We assume that the airframe has the following characteristics and constraints,
  • The structure of the quadrotor UAV is symmetric;
  • Friction between propeller and motor spindle is negligible;
  • The stator magnetic field speed and rotor speed of the motors are infinitely close to each other, such that the slip rate is 00;
  • The quadrotor and propeller structures are rigid, with a uniform mass distribution and the Geometric center is the Mass Center;
  • The Euler angle is bounded, i.e., $-\pi/2<\phi<\pi/2,-\pi/2<\theta<\pi/2,-\pi<\psi<\pi$.
Based on the above assumptions, we combine the dynamics principles to establish a rigid body model for attitude control of quadrotor UAV. It should be noted that all UAV coordinate system transformations are obtained by rotation synthesis with respect to the fixed coordinate system, the rotation matrix for the conversion of the body frame to the earth frame with the following equation,
$\begin{array}[]{l}R_{b}^{e}=\left(R_{e}^{b}\right)^{-1}=\left(R_{e}^{b}\right)^{T}=\\ \\ \left[\begin{array}[]{cll}\cos\theta\cos\psi&\sin\theta\cos\psi\sin\phi&\sin\theta\cos\psi\sin\phi\\ &-\sin\psi\cos\phi&+\sin\psi\cos\phi\\ \cos\theta\sin\psi&\sin\theta\sin\psi\sin\phi&\sin\theta\sin\psi\sin\phi\\ -\sin\theta&+\cos\psi\cos\phi&-\cos\psi\cos\phi\\ &\cos\theta\sin\phi&\cos\theta\cos\phi\end{array}\right]\end{array}$ (27)
Four inputs can be controlled based on the torque of the four motors of the UAV, including pitch control quantity $U_{1}$, roll control quantity $U_{2}$, yaw control quantity $U_{3}$, and altitude control quantity $U_{4}$,
$\left\{\begin{array}[]{l}U_{1}(t)=C_{T}(t)\left(\omega_{1}(t)^{2}+\omega_{2}(t)^{2}+\omega_{3}(t)^{2}+\omega_{4}(t)^{2}\right)\\ U_{2}(t)=C_{T}(t)\left(-\omega_{2}(t)^{2}+\omega_{4}(t)^{2}\right)\\ U_{3}(t)=C_{T}(t)\left(-\omega_{1}(t)^{2}+\omega_{3}(t)^{2}\right)\\ U_{4}(t)=C_{M}(t)\left(-\omega_{1}(t)^{2}+\omega_{2}(t)^{2}-\omega_{3}(t)^{2}+\omega_{4}(t)^{2}\right)\end{array}\right.$ (28)